OpenStudy (jenny12):

Iodine-131 has a half life of about 8 days. About how much is left from a 100 gram sample after 16 days?

OpenStudy (anonymous):

\[x/100 = (1/2)^{16/8}\] x/100 = (1/2)^2 x/100 = 1/4 x = 25

OpenStudy (jenny12):

Thank you.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

There are 2 half-lives involved here. The first half life will cut the sample down to 50g. The second half life will cut it down to 25 grams. It is half of what was left at each half-life period. Hope this helps. I had to do one of these equations in my Calc 3 class.

OpenStudy (jenny12):

So would that be 25 left?

OpenStudy (anonymous):

Yes. If you had a 24 minute period, then you would have 12.5g left, and so on.

OpenStudy (jenny12):

K, thank you.